A sample of hydrogen gas has a density of _____ g/L at a pressure of 0.581 atm and a temperature of 39 degrees Celsius. Assume ideal behavior.
Answer #1

pressure = P = 0.581 atm

temperature = 39 + 273 = 312 K

P V   = n R T
P V = W R T / M

P = dRT / M

d = P M / R T
d = 0.581 x 2 / 0.0821 x 312

d = 0.0454 g / L

density = 0.0454 g / L